# Dual-Domain Homogeneous Fusion with Cross-Modal Mamba and Progressive Decoder for 3D Object Detection

### v1

@@ Abstract
Fusing LiDAR point cloud features and image features in a homogeneous BEV space has been widely adopted for 3D object detection in autonomous driving. However, such methods are limited by the excessive compression of multi-modal features. While some works explore feature fusion in dense voxel spaces, they suffer from high computational costs and inefficiencies in query generation. To address these limitations, we propose a Dual-Domain Homogeneous Fusion network (DDHFusion), which leverages the complementary advantages of both BEV and voxel domains while mitigating their respective drawbacks. Specifically, we first transform image features into BEV and sparse voxel spaces using LSS and our proposed semantic-aware feature sampling module which can significantly reduces computational overhead by filtering unimportant voxels. For feature encoding, we design two networks for BEV and voxel feature fusion, incorporating novel cross-modal voxel and BEV Mamba blocks to resolve feature misalignment and enable efficient yet comprehensive scene perception. The output voxel features are injected into the BEV space to compensate for the loss of 3D details caused by height compression. For feature decoding, a progressive query generation module is implemented in the BEV domain to alleviate false negatives during query selection caused by feature compression and small object sizes. Finally, a progressive decoder can sequentially aggregate not only context-rich BEV features but also geometry-aware voxel features, ensuring more precise confidence prediction and bounding box regression. On the NuScenes dataset, DDHfusion achieves state-of-the-art performance, and further experiments demonstrate its superiority over other homogeneous fusion methods.
@@ I Introduction
Autonomous driving has emerged as a highly active research field, advancing rapidly with the integration of deep learning models. Among its critical technologies, 3D object detection plays a pivotal role in enabling safe and efficient driving planning. In current autonomous systems, LiDAR and cameras serve as essential sensors for environmental perception. LiDAR utilizes laser echo technology to capture precise 3D environmental information in the form of point clouds, while cameras generate images rich in texture information by capturing visible light. By harnessing the complementary strengths of LiDAR and cameras, fusion-based object detectors have consistently outperformed their single-sensor counterparts in terms of accuracy and robustness. However, multi-sensor feature fusion is non-trivial due to the inherent heterogeneity between camera and LiDAR data. Many existing works [1, 2, 3, 4, 5, 6] attempt to enrich raw point clouds or voxels with image features. However, this approach suffers from semantic information loss during camera-to-LiDAR projection, as only a small fraction of image features are preserved, leading to degraded detection performance in regions with fewer points. Recent studies propose homogeneous fusion strategies [7, 8, 9, 10, 11, 12] that actively transform image features into a shared representation space with LiDAR features. These methods leverage symmetric network architectures to fully exploit multi-modal complementarity while mitigating performance degradation caused by single-sensor failures.
@@
A popular method [7, 8, 9], as depicted in Fig. 1(a), employs Lift-Splat-Shoot (LSS) to project image features onto a unified BEV plane for concatenation with LiDAR features. This approach is highly efficient as mature techniques from 2D detection tasks can be easily adopted for BEV feature fusion and query generation. For the latter, the sample center-based label assignment and Top-K selection strategies are widely applied. However, it also has two limitations: 1) Information Compression: The compression of geometric and texture information along the height dimension results in blurred BEV features, which further leads to missed detections and reduced regression accuracy. As shown in Fig. 1(b), some works [10, 11] introduce multi-view encoder-decoder structures to address this. However, the additional transformation of LiDAR features back to the frontal view significantly disrupts 3D spatial relationships, still incurring semantic ambiguity. 2) Feature Misalignment: Inaccurate depth estimation in LSS or calibration errors between sensors may lead to spatial misalignment of multi-modal BEV features. Extra alignment techniques are required before feeding the features into the BEV backbone.
@@
As shown in Fig. 1(c), fusion in a unified voxel space has been explored to retain 3D geometric details and leverage higher-dimensional feature interactions for enhanced spatial perception while avoiding information compression. While theoretically superior to BEV-based fusion, this paradigm faces several challenges: 1) Computational Burden: Fusing features represented as dense voxels is highly memory-intensive and computationally expensive, making it difficult to deploy in large-scale driving scenarios. 2) Query Generation Difficulty: According to [13], compared to the BEV domain, the more severe foreground-background imbalance and higher dimensionality in the voxel domain complicate label assignment and input-dependent query generation [14]. [12] adopts randomly initialized 3D learnable queries combined with local 3D deformable attention in decoders, which slows down training convergence and limits the model’s ability to localize potential instances. 3) Feature Misalignment: The feature misalignment issue becomes even more challenging in the voxel domain. Computational constraints force these methods to rely on 3D convolutions for local feature fusion within limited receptive fields, which hinders feature alignment over larger regions.
@@
In this work, we propose a novel detection framework—the Dual-Domain Homogeneous Fusion network (DDHFusion), which overcomes the limitations of both BEV and voxel fusion domains while leveraging their respective strengths. As shown in Fig. 1(d), the BEV domain is responsible for query generation, while the voxel domain provides geometry-aware features. For clarity, the proposed method is divided into three steps: 1) Feature Transformation: First, we transform image features into sparse voxel and BEV spaces using the semantic-aware feature sampling (SAFS) and LSS modules. Compared to [12], the sparse voxel representation significantly reduces computational overhead. 2) Feature Encoding: During encoding, these multi-representation image features interact with LiDAR ones through homogeneous voxel and BEV fusion (HVF and HBF) networks. Within these networks, we introduce cross-modal voxel and BEV mamba blocks (CV-Mamba and CB-Mamba) to resolve feature misalignment and enable efficient global scene perception. 3) Feature Decoding: For decoding, we propose a progressive query generation (PQG) module and a progressive decoder (PD). The former activates BEV features with high-response queries, generating potential hard queries in the second stage to mitigate false negatives caused by feature compression and small object sizes. The latter enhances the conventional BEV decoder by integrating fine-grained 3D structural features around targets with our proposed multi-modal voxel feature mixing (MMVFM) module.
@@
    We introduce DDHFusion for 3D object detection that achieves state-of-the-art performance on the NuScenes dataset, outperforming other homogeneous fusion approaches while maintaining efficient inference.
@@
    To the best of our knowledge, we are the first to extend Mamba into multi-modal 3D detection. The proposed CV-Mamba and CB-Mamba blocks effectively facilitate feature fusion across global voxel or BEV domains.
@@
    We propose the PQG and PQ modules for feature decoding. The former generates easy and difficult queries in stages to minimize false negatives, while the latter adaptively integrates dual-domain features to enhance the accuracy of each query’s regression.
@@
In this section, we overview homogeneous fusion methods in the BEV and voxel domains, followed by recent advancements in state space models.
@@ II-A Homogeneous Fusion in BEV Domain
Fusion in a shared BEV space is a prominent approach for 3D object detection in autonomous driving. A critical step in this process is the transformation of image features into the BEV space. Early work [15] proposed an auto-calibration method: image features are projected onto predefined dense 3D voxel centers, adjusted by regressing offsets, and assigned to corresponding voxels. These features are then compressed along the height dimension to generate BEV features. Building on this, [16] introduced range attention to refine the distribution of intermediate voxel features along the range dimension. Recently, LSS [17] has become widely adopted for view transformation. It projects features into a 3D frustum based on depth estimation and then onto the BEV place with pooling operation. However, inaccuracies in depth estimation often lead to misalignment between LiDAR and image BEV views. To address this, [18] leverages contrastive learning for feature alignment, while [19] and [20] utilize semantic-guided optical flow estimation or mutual deformable attention for explicit alignment. Additionally, [21] and [20] enhance depth estimation accuracy through edge-aware LiDAR depth maps or local alignment operations.
Unlike the local fusion operations in the aforementioned methods, our approach establishes global relationships both within and across modalities, enabling effective feature alignment and more comprehensive scene perception.
@@ II-B Homogeneous Fusion in Voxel Domain
Compared to BEV homogeneous fusion, transforming image features into the voxel domain and fusing them with LiDAR voxels avoids the information loss caused by height compression. Additionally, introducing multi-modal voxel features during the decoding process leads to better regression accuracy. [22] and [23] unproject image features from each pixel onto virtual points with depth completion and then convert them into voxels, applying voxel pooling to fuse fine-grained multi-modal features. [24] generates virtual points using nearest neighbor matching around the 2D instance mask on the image. [25] improves this approach with k-nearest neighbor matching and introduces gated modality-aware convolution to fuse semantic and geometric features of both camera and LiDAR at different granularities. While these virtual-point-based methods largely address the sparsity issue of LiDAR point clouds, they require additional depth completion or high-resolution instance segmentation networks. It leads to significant training costs, and the labels required for these tasks are often difficult to obtain. In contrast, [12] bypasses the need for extra networks by adopting an LSS-like algorithm. It first assigns image features to predefined dense voxels and then weights them with category-based depth maps. [26] eliminates the depth-weighting step and introduces additional gating mechanisms to enable flexible fusion. However, dense voxel representations incur significant computational overhead. In recent work, [27] only utilize element-wise fusion in the intermediate voxel space during LSS, with subsequent feature propagation occurring in the BEV domain.
@@
In our work, we adopt sparse voxel representation to conserve computational resources. Additionally, our Mamba-based network, HVF, facilitates multi-granularity global feature fusion.
II-C State Space Model
@@
State Space Models (SSMs) [28, 29, 30] describe dynamic systems by representing the system’s internal state at each moment. With advancements in deep learning, SSMs have been integrated into neural networks to achieve more efficient and flexible inference. [31] effectively captures long-term dependencies in sequences by introducing higher-order polynomial projection operators, compressing historical information into a low-dimensional state space. S4 [28], a structured state space model based on [29], enhances computational efficiency and modeling capacity through structured state transition matrices. Building on S4, [32] introduces input-specific state transition mechanisms, significantly boosting both expressiveness and efficiency. Due to its linear complexity, Mamba has been widely adopted in computer vision. For instance, [33] unfolds images into 1D sequences and proposes a bidirectional Mamba module for comprehensive feature learning. [34] extends this approach to four-directional scanning, uncovering richer spatial relationships in images and providing a theoretical analysis of Mamba’s linear attention. Mamba has also been applied to 3D tasks such as point cloud classification [35], 3D object detection [36, 37, 38], and semantic completion [39]. For example, [36] employs a group-free voxel Mamba module to expand the receptive field across the entire scene, while [39] integrates local adaptive reordering into the Mamba module to enhance local information extraction.
@@
Different from these single-modality architectures for 3D tasks, we extend Mamba to multi-modal ones, further unlocking its potential.
@@ III Method
The pipeline of DDHFusion is shown in Fig. 2. First, multi-view images and LiDAR point clouds are passed through their respective networks for feature extraction. These features are then transformed into sparse voxels $V_{I}$, $V_{L}$ and BEV representations $B_{I}$,$B_{L}$. Then, two homogeneous fusion networks are introduced to perform feature fusion within their respective domains. The outputs $V_{I}^{{}^{\prime}}$ and $V_{L}^{{}^{\prime}}$ are passed into the BEV fusion network after sparse height compression [40]. After processing through the BEV backbone, the resulting BEV features are used for progressive query generation. These queries are then input into the progressive decoder including both BEV and voxel decoders for classification and box regression. In this section, we provide a detailed introduction to DDHFusion from the following aspects: semantic-aware feature sampling, voxel homogeneous network, BEV homogeneous network, progressive query generation, and progressive decoder.
III-A Semantic-Aware Feature Sampling
@@
Describing the entire driving scene using dense voxels is highly redundant. First, a significant portion of voxels corresponds to empty space, where no objects are present. Second, in autonomous driving scenarios, foreground objects typically occupy only a small fraction of the overall scene. This redundancy not only consumes substantial computational resources but also diverts the model’s information extraction capacity to large, irrelevant areas, thereby hindering feature learning.
@@
Drawing inspiration from [41], we propose the SAFS module depicted in Fig. 3 to selectively generate important image voxels. First, the image features $F_{I}$ from the image backbone are passed through a convolution block to generate the depth map. Similar to LSS, these depths are represented by discrete bins. Another branch predicts the semantic segmentation mask. During training, we project points within ground truth boxes onto the image plane to generate segmentation labels for sparse supervision.
@@
Figure 2: Overview of DDHFusion. It begins by extracting features from multi-view images and LiDAR point cloud and transforms both into the voxel and BEV domains. These features are then passed to two Mamba-based homogeneous fusion networks, which perform feature alignment and global perception. The resulting high-quality fused BEV feature $B_{out}$ is then fed into the progressive query generation module, which leverages the spatial relationship of easy queries to stimulate the generation of hard queries. Finally, all queries are passed into the progressive decoder, which abstracts dual-domain features around instances for accurate classification and box regression.
@@
Figure 3: Details of semantic-aware feature sampling (SAFS). We filter unimportant voxels by depth and semantic scores and transform the image feature to the sparse voxel domain.
@@
At the beginning of training or in some crowd scenarios, the number of $V_{c}^{{}^{\prime}}$ may be too large. Therefore, we set an upper limit $N$, and if the number exceeds it, the farthest point sampling is applied to retain only $N$ voxels. Finally, we assign image features to $V_{c}^{{}^{\prime}}$ and multiply them by the corresponding depth scores to obtain voxel features $V_{I}$.
III-B Homogeneous Voxel Fusion Network
In HVF, we employ a 3D U-Net architecture to fuse LiDAR voxels $V_{L}$ and image voxels $V_{I}$. As illustrated in Fig. 4, the network consists of two parallel branches that process voxel features separately, with feature fusion performed at each scale. Due to the uncertainty of depth estimation and calibration error, $V_{L}$ and $V_{I}$ are often misaligned spatially. Previous works [12, 22, 25] only fuse voxel features by concatenation or local gating mechanisms within the same position or a $3\times 3\times 3$ neighborhood. However, they fail to resolve the misalignment problem and limit the network’s ability to comprehensively perceive multi-modal information. Inspired by the recent success of Mamba in LiDAR-based 3D object detection [36, 37], we propose intra-modal and cross-modal voxel Mamba (IV-Mamba and CV-Mamba) modules for feature fusion at each scale. In IV-Mamba, we rearrange the voxels into a 1D sequence based on the 3D Hilbert curve according to [36]. The discrete SSM model is then used to process the entire sequence:

### v2

@@ Abstract
Fusing LiDAR and image features in a homogeneous BEV domain has become popular for 3D object detection in autonomous driving. However, this paradigm is constrained by the excessive feature compression. While some works explore dense voxel fusion to enable better feature interaction, they face high computational costs and challenges in query generation. Additionally, feature misalignment in both domains results in suboptimal detection accuracy. To address these limitations, we propose a Dual-Domain Homogeneous Fusion network (DDHFusion), which leverages the complementarily of both BEV and voxel domains while mitigating their drawbacks. Specifically, we first transform image features into BEV and sparse voxel representations using lift-splat-shot and our proposed Semantic-Aware Feature Sampling (SAFS) module. The latter significantly reduces computational overhead by discarding unimportant voxels. Next, we introduce Homogeneous Voxel and BEV Fusion (HVF and HBF) networks for multi-modal fusion within respective domains. They are equipped with novel cross-modal Mamba blocks to resolve feature misalignment and enable comprehensive scene perception. The output voxel features are injected into the BEV space to compensate for the information loss caused by direct height compression. During query selection, the Progressive Query Generation (PQG) mechanism is implemented in the BEV domain to reduce false negatives caused by feature compression. Furthermore, we propose a Progressive Decoder (QD) that sequentially aggregates not only context-rich BEV features but also geometry-aware voxel features with deformable attention and the Multi-Modal Voxel Feature Mixing (MMVFM) block for precise classification and box regression. On the NuScenes dataset, DDHFusion achieves state-of-the-art performance, and extensive experiments demonstrate its superiority over other representative homogeneous fusion methods.
@@ I Introduction
Autonomous driving has emerged as a prominent research frontier, advancing rapidly with the integration of deep learning models. Among its technologies, 3D object detection plays a pivotal role in enabling safe and efficient driving planning. In current autonomous systems, LiDAR and cameras serve as essential sensors for environmental perception. LiDAR utilizes laser echo technology to capture precise 3D spatial information in the form of point clouds, while cameras generate images rich in texture details by capturing visible light. By harnessing the complementary strengths of LiDAR and cameras, fusion-based object detectors have consistently outperformed their single-sensor counterparts in terms of accuracy and robustness. However, multi-sensor feature fusion is non-trivial due to the inherent heterogeneity between camera and LiDAR data. Many existing works [1, 2, 3, 4, 5, 6] simply enrich LiDAR points or voxels with image features by camera-to-LiDAR projection. However, this approach suffers from the loss of semantic information as only a small fraction of image features are preserved, leading to degraded detection performance in areas with sparse point distributions. Recent studies propose homogeneous fusion strategies [7, 8, 9, 10, 11, 12] that actively transform image features into a shared space with LiDAR features. These approaches employ symmetric network architectures to exploit the combined advantages of multi-modal inputs, while simultaneously reducing the performance degradation that arises from single-sensor failures
@@
A popular paradigm [7, 8, 9], as depicted in Fig. 1(a), employs Lift-Splat-Shoot (LSS) to project image features onto a unified BEV plane, where they are concatenated with LiDAR features. Its framework is highly efficient as mature techniques from 2D detection tasks can be easily adopted for BEV feature fusion and query generation. For the latter, the simple center-based label assignment and Top-K selection strategies are widely applied. However, there remain two limitations: 1) Information Compression: Collapsing geometric and texture information along the height dimension results in blurred features, which further leads to missed detections and reduced regression accuracy. As shown in Fig. 1(b), some works [10, 11] introduce multi-view encoder-decoder structures to address this. However, the additional transformation of LiDAR features to the frontal view significantly disrupts 3D spatial relationships, still incurring feature ambiguity. 2) Feature Misalignment: Inaccurate depth estimation in LSS between sensors may lead to spatial deviations of multi-modal BEV features. Extra alignment techniques are required before feeding the features into the BEV backbone.
@@
As shown in Fig. 1(c), fusion in a unified dense voxel space [12] has been explored to and leverage higher-dimensional feature interactions while avoiding information compression. While theoretically superior to BEV-based fusion, it faces several challenges: 1) Computational Burden: Fusing features represented as dense voxels is highly memory-intensive and computationally expensive, making it difficult to deploy in large-scale driving scenarios. 2) Query Generation Difficulty: According to [13], compared to the BEV domain, the more severe foreground-background imbalance and higher dimensionality in the dense voxel domain complicate label assignment and input-dependent query generation [14]. [12] adopts randomly initialized 3D learnable queries combined with local 3D deformable attention in decoders, which slows down training convergence and limits the model’s ability to localize potential instances. 3) Feature Misalignment: During feature transformation, the depth estimation, similar to that in LSS, also leads to feature misalignment. Worse still, computational constraints compel the network to rely on 3D convolutions for local fusion within limited receptive fields, which hinders feature alignment over larger regions.
@@
In this work, we propose a novel detection framework—the Dual-Domain Homogeneous Fusion network (DDHFusion), which overcomes the limitations of BEV and voxel fusion domains while leveraging their respective strengths. As shown in Fig. 1(d), the BEV domain is responsible for query generation, while the voxel domain provides geometry-aware features during feature encoding and decoding. For clarity, the proposed DDHFusion is divided into three steps: 1) Feature Transformation: First, we transform image features into sparse voxel and BEV spaces using the semantic-aware feature sampling (SAFS) and LSS modules. Compared to [12], the sparse voxel representation significantly reduces computational overhead. 2) Feature Encoding: During encoding, these multi-representation image features interact with LiDAR ones through homogeneous voxel and BEV fusion (HVF and HBF) networks. Within these networks, we meticulously designs cross-modal voxel and BEV mamba blocks (CV-Mamba and CB-Mamba) to address feature misalignment and enable efficient global scene perception. 3) Feature Decoding: For decoding, we propose a progressive query generation (PQG) module and a progressive decoder (PD). The former activates BEV features with the spatial relationship of high-response queries, generating potential hard queries in the second stage to mitigate false negatives caused by feature compression and small object sizes. The latter level up the conventional BEV decoder by integrating our proposed multi-modal voxel feature mixing (MMVFM) module which can aggregate fine-grained 3D structural information around targets for precise box regression.
@@
    We present DDHFusion, a novel framework for 3D object detection that achieves state-of-the-art performance on the NuScenes dataset. It surpasses other homogeneous fusion methods in accuracy while maintaining efficient inference.
@@
    To the best of our knowledge, we are the first to extend Mamba into the multi-modal 3D detection task. The proposed CV-Mamba and CB-Mamba blocks effectively facilitate feature fusion across global voxel and BEV domains.
@@
    We propose PQG and PQ modules for feature decoding. The former generates easy and hard queries in stages to minimize false negatives, while the latter adaptively integrates dual-domain features to enhance regression accuracy of each query.
@@
In this section, we overview homogeneous fusion methods in BEV and voxel domains. Additionally, since Mamba is utilized in this work, we also introduce recent advancements in state space models.
@@ II-A Homogeneous Fusion in BEV Domain
The introduction of homogeneous BEV fusion marks a significant milestone in the evolution of perception techniques for autonomous driving. A critical step in this method is the transformation of image features into the BEV space. The early study [15] proposes an auto-calibration mechanism: image features are assigned to predefined voxel centers through 2D-3D projection and offset adjustment. These features are then compressed along the height dimension to generate BEV features. Building on this, [16] further introduces range attention to refine the distribution of intermediate voxel features along the range dimension. Recently, LSS [17] has become widely adopted for view transformation. It projects features into a 3D frustum based on depth estimation and then onto the BEV plane with a pooling operation. However, inaccuracies in depth estimation often lead to misalignment between LiDAR and image BEV features. To address this, [18] leverages contrastive learning for feature alignment, while [19] and [20] utilize semantic-guided optical flow estimation or mutual deformable attention for explicit alignment. Additionally, [21] and [20] attempt to enhance depth estimation accuracy through edge-aware LiDAR depth maps or local graph alignment in the frontal view.
Unlike the local fusion operations in the aforementioned methods, our approach establishes global relationships both within and across modalities, enabling effective feature alignment and a more comprehensive understanding of the driving scene.
@@ II-B Homogeneous Fusion in Voxel Domain
Compared to BEV homogeneous fusion, transforming image features into the voxel domain and fusing them with LiDAR voxels avoids the information loss caused by height compression. Additionally, introducing multi-modal voxel features during the decoding process leads to better regression accuracy. [22] and [23] unproject image features from each pixel onto virtual points with depth completion and then convert them into voxels, applying voxel pooling to fuse fine-grained multi-modal features. [24] generates virtual points using nearest neighbor matching within 2D instance masks on the image. [25] improves this approach with k-nearest neighbor matching and introduces gated modality-aware convolution to fuse semantic and geometric features of both camera and LiDAR at different granularities. While these virtual-point-based methods largely address the sparsity issue of LiDAR point clouds, they require extra depth completion or high-resolution instance segmentation networks. This leads to significant training costs and labels required for these tasks are often difficult to obtain. In contrast, [12] bypasses the need for extra networks by adopting an LSS-like algorithm. It assigns image features to predefined dense voxels and then weights them with category-based depth maps. [26] eliminates the depth-weighting step and introduces additional gating mechanisms to enable flexible fusion. However, dense voxel representations incur significant computational overhead. In recent work, [27] only utilizes element-wise fusion in the intermediate voxel space during LSS, with subsequent feature propagation occurring in the BEV domain.
@@
In our work, we adopt sparse voxel representation to conserve computational resources. Additionally, our Mamba-based network—HVF, facilitates multi-granularity global feature fusion.
II-C State Space Models
@@
State Space Models (SSMs) [28, 29, 30] describe dynamic systems by representing the system’s internal state at each moment. With advancements in deep learning, SSMs have been integrated into neural networks to achieve more efficient and flexible inference. [31] effectively captures long-term dependencies in sequences by introducing higher-order polynomial projection operators, compressing historical information into a low-dimensional state space. S4 [28], a structured state space model based on [29], enhances computational efficiency and modeling capacity through structured state transition matrices. Building on S4, [32] introduces an input-specific state transition mechanism–Mamba, significantly boosting both expressiveness and efficiency. Due to its linear complexity and modeling strength comparable to Transformer, many studies have explored its application in computer vision. For instance, [33] unfolds images into 1D sequences and proposes a bidirectional Mamba module for comprehensive feature learning. [34] extends this approach to four-directional scanning, uncovering richer spatial relationships in images. It also reveals the connection between Mamba and linear attention through theoretical analysis, providing deeper insights into its underlying mechanisms. Mamba has also been applied to 3D tasks such as point cloud classification [35], 3D object detection [36, 37, 38] and semantic completion [39]. For example, [36] employs a group-free voxel Mamba module to expand the receptive field to the entire scene, while [39] integrates local adaptive reordering into the Mamba module to enhance local information extraction.
@@
Different from these single-modality architectures for 3D tasks, we adapt Mamba to the multi-modal framework, further unlocking its potential of versatile applications.
@@ III-A Overview
The pipeline of DDHFusion is shown in Fig. 2. Initially, multi-view images and LiDAR points are passed through their respective backbones for feature extraction. The LiDAR voxels $V_{L}$ are converted into $B_{L}$ through height compression, while image features are transformed into sparse voxels $V_{I}$, BEV map $B_{I}$ via LSS and SAFS modules. Then, we meticulously design the homogeneous voxel fusion network, where intra-modal voxel Mamba and cross-modal voxel Mamba (IV-Mamba and CV-Mamba) blocks are alternately employed through merge-and-split operations to interact $V_{I}$ and $V_{L}$. The resulting voxels $V_{I}^{{}^{\prime}}$ and $V_{L}^{{}^{\prime}}$ are fed into the homogeneous BEV fusion network and converted into $B_{L}^{\prime}$, $B_{L}^{\prime}$ by sparse height compression [40]. The multi-modal BEV features $B_{L}$, $B_{I}$, $B_{L}^{\prime}$ and $B_{I}^{\prime}$ are fused using intra-modal and cross-modal BEV Mamba blocks (IB-Mamba and IB-Mamba) and the result is then processed through the BEV backbone for deep feature extraction. The resulting $B_{out}$ is used for progressive query generation. Finally, both easy queries $Q_{e}$ and hard queries $Q_{h}$ are concatenated and input into the progressive decoder which includes both BEV and voxel decoders for classification and box regression. In the remainder of this section, we provide a detailed introduction to DDHFusion from the following aspects: semantic-aware feature sampling, voxel homogeneous network, BEV homogeneous network, progressive query generation, and progressive decoder.
III-B Semantic-Aware Feature Sampling
@@
Describing the entire driving scene using dense voxels is highly redundant. First, a substantial portion of voxels corresponds to empty space, where no objects exist. Second, in autonomous driving scenarios, foreground objects typically occupy only a small fraction of the overall scene. This redundancy not only consumes considerable computational resources but also diverts the network’s modeling capacity to large, irrelevant areas, thereby hindering feature learning.
@@
Drawing inspiration from [41], we propose the SAFS module depicted in Fig. 3 to selectively generate important image voxels. First, the image feature $F_{I}$ from the image backbone are passed through a convolution block to generate the depth map. Similar to LSS, these depths are represented by discrete bins. Another branch predicts the semantic segmentation mask. During training, we project points within ground truth boxes onto the image plane to generate segmentation labels for sparse supervision.
@@
Figure 2: Overview of DDHFusion. It begins by extracting features from multi-view images and LiDAR points and transforms both into the voxel and BEV domains. These features are then passed to two Mamba-based homogeneous fusion networks, which perform feature alignment and global perception. The resulting high-quality fused BEV feature $B_{out}$ is then fed into the progressive query generation module, which leverages the spatial relationship of easy queries to stimulate the generation of hard queries. Finally, all queries are passed into the progressive decoder, which abstracts dual-domain features around instances for accurate classification and box regression.
@@
Figure 3: Details of semantic-aware feature sampling (SAFS). We filter the coordinates of unimportant voxels by depth and semantic scores and transform the image feature to the sparse voxel domain.
@@
At the beginning of training or in some crowd scenarios, the number of $V_{c}^{{}^{\prime}}$ may be too large. Therefore, we set an upper limit $N$. If the number exceeds it, the farthest point sampling is applied to retain only $N$ voxels. Finally, we assign image features to $V_{c}^{{}^{\prime}}$ and multiply them by the corresponding depth scores to obtain voxel features $V_{I}$.
III-C Homogeneous Voxel Fusion Network
As illustrated in Fig. 4, the HVF consists of two parallel branches that process voxel features separately, with feature fusion performed at each scale. Existing works [12, 22, 25] fuse multi-modal voxel features through element-wise addition or local gating mechanisms limited to a small neighborhood. However, due to the uncertainty of depth estimation during feature transformation, the corresponding latent representations in $V_{L}$ and $V_{I}$ are often spatially deviated. With limited search ranges, these local fusion operations fail to effectively address this misalignment problem. Inspired by the recent success of Mamba in LiDAR-based 3D object detection [36, 37], we propose IV-Mamba and CV-Mamba modules in the HVF for global feature interaction, which can better correlate severely misaligned multi-modal information and enable holistic understanding of the scene.
In IV-Mamba, following [36], we rearrange the voxels into a 1D sequence in the order of the 3D Hilbert curve for its preservation of spatial proximity. The discrete SSM model is then used to process the entire sequence: